OpenStudy (istim):

Determine the points on each curve that correspond to the given slope of the tangent. y=(2x-1)(-4+x^2), m=3

OpenStudy (istim):

My final answers are (4/3,-3.703) and (-1,9). The correct answers are (1.53,-3.42) and (-1.20,8.70)

OpenStudy (dumbcow):

don't understand the question are there 2 functions?

OpenStudy (istim):

Only one function, but the slope is 3.

OpenStudy (dumbcow):

ok so set dy/dx = 3 dy/dx = 2(x^2-4) + 2x(2x-1) = 3 --> 6x^2 -2x -11 = 0 use quadratic formula to get x_coordinates

OpenStudy (istim):

That's what I think did, but I got the answer wrong.

OpenStudy (dumbcow):

check your computation of quadratic formula...the above function gives correct x-values...-1.197, 1.53

OpenStudy (istim):

Well. Here's what I did and results: -Found dy/dx=6x^2-2x-8. -Factored out 2 so 0=2(3x^2-x-4) -Oh, didn't do quadratics. Slip of the mind, sorry. -Factored the equation so x=4/3 amd x=-1.

OpenStudy (dumbcow):

oh i see...the only part that is wrong is you assumed it equaled 0 but RHS should be 3 6x^2 -2x -8 = 3 then factoring out the 2 doesn't really help

OpenStudy (istim):

Ok. I'll try that. Thanks. I'll Be off for 15 minutes.